9*(x+2/x+3)^2 - 6*(x+2/x+3)+1 = 0
Answer by stanbon(75887)

9*(x+2/x+3)^2 - 6*(x+2/x+3)+1 = 0
----------
Let w = (x+2/x+3)
9w^2 - 6w + 1 = 0
(3w -1)(3w - 1) = 0
w = 1/3
---------------
(x+2)/(x+3) = 1/3
Cross-multiply
x+2 = 3x+9
2x = -7
x = -7/2
==============
Cheers,
Stan H.